# Link Representation Learning for Probabilistic Travel Time Estimation

### v1

@@ Abstract
Travel time estimation is a crucial application in navigation apps and web mapping services. Current deterministic and probabilistic methods primarily focus on modeling individual trips, assuming independence among trips. However, in real-world scenarios, we often observe strong inter-trip correlations due to factors such as weather conditions, traffic management, and road works. In this paper, we propose to model trip-level link travel time using a Gaussian hierarchical model, which can characterize both inter-trip and intra-trip correlations. The joint distribution of travel time of multiple trips becomes a multivariate Gaussian parameterized by learnable link representations. To effectively use the sparse GPS trajectories, we also propose a data augmentation method based on trip sub-sampling, which allows for fine-grained gradient backpropagation in learning link representations. During inference, we estimate the probability distribution of the travel time of a queried trip conditional on the completed trips that are spatiotemporally adjacent. We refer to the overall framework as ProbTTE. We evaluate ProbTTE on two real-world GPS trajectory datasets, and the results demonstrate its superior performance compared to state-of-the-art deterministic and probabilistic baselines. Additionally, we find that the learned link representations align well with the physical geometry of the network, making them suitable as input for other applications.
@@
Probabilistic travel time estimation, representation learning, uncertainty quantification, low-rank parameterization
@@ I Introduction
Travel Time Estimation (TTE) focuses on predicting the time it takes a vehicle to traverse a specific path connecting two locations (i.e., origin and destination). Given the departure time, TTE is equivalent to predicting the estimated time of arrival (ETA). TTE (or ETA prediction) is a fundamental application in navigation services and intelligent transportation systems: for example, individual travelers can better schedule/plan their trips based on TTE, logistics and delivery service operators can design more efficient and robust routes by taking TTE into consideration, and transport agencies can leverage TTE for better operation and management.
@@
Given its importance, ETA prediction has become an important problem in machine learning and has gained considered attention in the literature. Most existing studies can be categorized into two groups based on whether path information is used to predict travel time [1]. The first is the origin-destination-based (OD-based) approach (see, e.g., [2], [3], [4], [5], [6], [7], [8]). The key concept of this approach involves searching for past trips that have origin-destination profiles similar to the queried trip. The obtained travel time values are then used to calculate the mean and variance of the queried travel time [3]. Factors such as travel distance, departure time, road attributes, and other additional information are further introduced to enrich the feature space of OD pairs [4], [6], [8]. The periodic patterns of traffic conditions can also be analyzed to account for the temporal dynamics [5], [7], [8]. The OD-based method relies solely on structured features such as origin, destination, and time of day as covariates, and the models typically exhibit low computational complexity. Nonetheless, since travel time is primarily influenced by the specific route taken by a vehicle, the prediction accuracy of these OD-based models is often compromised due to the omission of path information. The presence of high-quality GPS data and advances in sequence modeling methodologies inspire researchers to use comprehensive trip information for TTE, resulting in the development of route-focused approaches. The key concept of the route-based approach involves learning representation vectors for road segments [9], [10] or GPS points[11] in a data-driven way and then combining these vectors with deep neural networks to obtain estimates. Most route-based methods utilize recursive neural networks to model temporal dynamics (see, e.g., [12], [13], [14], [15]) and graph neural networks to characterize spatial information (see, e.g., [16], [17], [18]), [19]. In addition, attention mechanism [20], [21] and meta-learning [22], [23] are also incorporated into this category of methods.
Due to the importance of TTE in travel decision-making, such as when to leave and which path to take, recent research has focused on probabilistic TTE, which involves forecasting the probabilistic distribution of travel time on a specific path [24], [25]. These models generally assume that the travel time $\tau_{i}$ of a trip $i$ follows a Gaussian distribution $\tau_{i}\sim\mathcal{N}\left(f_{\mu}(X_{i}),f_{\sigma^{2}}(X_{i})\right)$, where $X_{i}$ is the input feature of trip $i$, and $f_{\mu}$ and $f_{\sigma^{2}}$ are two neural networks that produce the mean and variance of the distribution, respectively. To facilitate likelihood calculation, trips are assumed to be independent, i.e., $\operatorname{Cov}(\tau_{i},\tau_{j})=0,\forall i\neq j$. However, considering the spatiotemporal correlation in traffic conditions, the travel time for two trips could be correlated if both are influenced by unseen factors (e.g., incidents and weather conditions). Disregarding this information can result in less effective representation learning for road segments within the traffic network. A natural way to model such correlation is to design a random effects model for segment travel times (see e.g., [1]), and consider trip travel time as an aggregated observation of segment travel times together with trip-level random effects. However, in practice, this is challenging because of the large dimensionality of the problem: to model the spatial correlation of segment travel time, two large covariance matrices, for inter-trip segment travel time and intra-trip segment travel time, respectively, need to be estimated from the aggregated travel time observations. For a transportation network with $N$ segments, the covariance matrix is of size $N\times N$, which is computationally prohibitive for large networks. In addition, due to the sparse nature of GPS data, segment travel time for an individual trip often cannot be observed directly. Instead, the observed time interval between two consecutive GPS data points often reflects the aggregated travel time over multiple road segments. This also poses challenges in attributing the variation in total travel time on a path to each individual link on the path.
To effectively use GPS trajectories for learning trip travel time distributions, in this paper we propose the ProbTTE framework, which models the joint distribution of multiple trips as a multivariate Gaussian distribution. In particular, we design the covariance matrix to account for both inter-trip correlations and intra-trip correlations through learnable link/segment representations. In terms of training, we leverage the low-rank-plus-diagonal structure of the covariance matrix, and use negative log-likelihood as a loss function to update link representations. The link representation learned by ProbTTE also establishes a unified projection with a contraction property, which allows us to build efficient representations for trips through affine transformations of link representations. Additionally, a trip sub-sampling data augmentation method is developed to utilize the sparse GPS trajectories more efficiently for link representation learning. The main contributions of this paper are summarized as follows:
@@
  • We propose a multi-trip joint probabilistic distribution model ProbTTE for estimating travel times. Efficient low-rank parameterizations are introduced to learn inter-trip and intra-trip correlations among all segments in a transportation network.
@@
  • We propose a sub-sampling data augmentation approach to balance the samples and enhance the optimization efficiency in learning the link representation vectors. It facilitates specific optimization at the link level, enabling fine-grained modeling of link features.
@@
The remainder of this paper is organized as follows. Section II summarizes related work. Section III introduces key definitions and formulations. Section IV presents the key components of ProbTTE, including parameterization of the inter-trip and intra-trip link covariance matrices and efficient likelihood evaluation. We evaluate the proposed model on two real-world datasets in Section V. Finally, we conclude this study in Section VI.
@@ II-A Deterministic Travel Time Estimation
ProbTTE(ours) Route-based GPS Sequence Link Embedding Path-based-sum of Link Representation
@@ II-B Probabilistic Regression
In TTE, there are also some efforts to explore probabilistic modeling. Li et al. [25] proposed a deep generative model to learn the travel time distribution by conditioning on the real-time traffic. This model interprets the generation of travel time using a three-layer hierarchical probabilistic model, which captures dynamically changing real-time traffic conditions and static spatial features, and then generates estimation times based on an attention mechanism. In [24], Zhou et al. proposed to learn the local representations of road segments over a temporal attributed graph, by jointly exploiting the dynamic traffic conditions and the topology of the road networks. Then a distribution loss based on the negative log-likelihood (NLL) is developed to fulfill the purpose of travel time distribution estimation. Mao et al. [26] introduced GMDNet, a Graph-based Mixture Density Network, which harnesses the advantages of both graph neural networks and mixture density networks for estimating travel time distribution. They utilized the Expectation-Maximization (EM) framework to enhance stability during training. In addition, tensor-based methods have also been used to construct probabilistic estimation models for travel time.
@@ III-B Definition: Links and Trips
A link $l\in\mathcal{V}$ refers to a road segment. Each link in the road network has a unique index. In this work, we follow OpenStreetMap (OSM) [29] to define links. The trajectory of a trip is an ordered sequence of time-stamped GPS records. For a trip with $k$ GPS data points, we denote its trajectory by $T_{g}=\left\{(l_{1},c_{1}),\ldots,(l_{k},c_{k})\right\}$, where the time index $c$ is monotonically increasing. The total travel time of the trip is $\tau=c_{k}-c_{1}$. Note that it is possible that (1) more than one GPS data points are located on the same link, and (2) a traversed link is not captured if no GPS records are registered. The link index sequence, after completing the entire sequence and removing duplicates, is used as model input, denoted by ${T}=\{l_{1},...,l_{k}\}$. In other words, a trip $T$ can be regarded as an ordered set of traveled links.
@@ IV Methodology
IV-A Overview of ProbTTE
The overall architecture of ProbTTE is shown in Figure 1. A unique property of ProbTTE is that the correlation among multiple trips is explicitly modeled, and the travel times for multiple trips are jointly modeled as a multivariate Gaussian distribution. Each trip is represented by aggregating the low-rank representation of all those links that it covers. A neural network-based mapping function is constructed to learn the mean travel time based on the trip representation vectors, while the covariance is formed using the inner product of these trip representation vectors for simultaneously modeling inter- and intra-trip correlations. Additionally, trip sub-sampling is used to augment the training data, balancing the samples and allowing gradients to differentially impact the representation at the link level within the same trip. Finally, historical trips in spatiotemporal proximity to the query trip are introduced to obtain the travel time distribution of a queried trip conditional on those observed trips. The final estimations are obtained by random sampling the conditional distribution. We explain the details of each component in the following subsections.
@@ IV-B Parameterizing Link Travel Time Distribution
In this study, we focus on modeling trip travel time over a specific time horizon (e.g., 8:00–9:00 AM) of weekdays. In doing so, we restrict our scope to only model the variations in link travel time resulting from external factors rather than traffic demand. Given that the historical data spans several days, we employ a hierarchical model to describe both daily random fluctuations and random effects at the trip level. In particular, we model the travel time $t_{l,i,q}$ for link $l$, day $i$, and trip $q$ as
@@
where $\mu_{l}$ represents the overall mean travel time for link $l$, which can be produced by link representation and other available covariates such as time of day and day of week, $\eta_{l,i}$ accounts for day-specific deviations as a result of the impacts of global unobserved factors that affects all trips in a day but are not used in modeling $\mu_{l}$, such as weather conditions, road works and public holidays; and $\epsilon_{l,i,q}$ captures the trip-specific error (e.g., differences in driver/vehicle profile, or short-duration delays due to traffic incidents affecting multiple nearby links). This specification is similar to [1], where day-specific and trip-specific random effects are introduced to capture travel time variations.
@@
$\Sigma_{d}=LL^{\top},$ (4)
where $L\in\mathbb{R}^{|\mathcal{V}|\times r_{L}}$ and $r_{L}\ll|\mathcal{V}|$. With this assumption, $\Sigma_{d}$ is positive semi-definite and we can consider each row of $L$ as a representation vector of a link/segment. We adopt a multilayer perceptron $f_{\mu}(\cdot)$ to learn the global mean of travel time based on the representation matrix $L$
$\bm{\mu}=L\bm{w}_{\mu},$ (5)
where $\bm{w}_{\mu}\in\mathbb{R}^{r_{L}}$ is a parameter vector to be learned. Combining Eqs. (5) and (4), we can specify the joint Gaussian distribution based on the link representation matrix $L$. It should noted that the modeling of $\bm{\mu}$ is not limited to the segment feature $L$. One can also introduce other features that are available, such as driver/vehicle profile, time of day, day of week, public holiday information, and weather conditions. This will create a more comprehensive mean process. For example, Ref. [30] introduces representation learning for rare temporal conditions such as events and holidays.
Figure 2: Graph node contraction.
In addition to computational efficiency, the proposed specification also provides excellent contraction properties and facilitates the aggregation of unimportant nodes in $\mathcal{V}$. For example, the contractions of links $\{1,2\}$ and $\{8,9,10\}$ in Figure 2 can be achieved by introducing a $7\times 10$ binary mapping matrix
@@
Given the affine property of the Gaussian distribution, the mean travel time in the simplified network is $\bm{\mu}^{\prime}=M\bm{\mu}=ML\bm{w}_{\mu}$ and the covariance matrix becomes $\Sigma_{d}^{\prime}=ML(ML)^{\top}$. This corresponds to having a new representation matrix $L^{\prime}=ML$. This affine property can further help us in deriving the distribution of trip travel time. We define the representation of a trip (as a row vector) to be the path-based-sum of link representation vectors, i.e., $ML$, with $M\in\{0,1\}^{1\times|\mathcal{V}|}$. The mean travel time on a path becomes $M(L\bm{w}_{\mu})=(ML)\bm{w}_{\mu}$ can be directly computed from the trip representation $ML$.
@@
$\Sigma_{p}=HH^{\top}+D,$ (7)
where $H\in\mathbb{R}^{|\mathcal{V}|\times r_{H}}$ with $r_{H}\ll|\mathcal{V}|$ and $D\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ is diagonal matrix with $D_{i,i}>0$ for $i=1,\ldots,|\mathcal{V}|$. In the neural network, we use a Softplus function nested with multilayer perception to model the diagonal entries $\text{diag}(D)=\log\left(1+\exp(f_{d}(H))\right)$ where $f_{d}(H)=H\bm{w}_{d}$ and $\bm{w}_{d}$ is a vector of length $r_{H}$. Similar to $L$, we can also perform contractions on the learned matrix $H$ and the vector $\operatorname{diag}(D)$.
IV-C Joint Distribution for Multiple Trips in a Day
A key challenge in learning the link representations is that trip-level link travel time $t_{l,i,q}$ is not observable in the raw GPS data. Instead, what we have is the final travel time for a trip. Let $\bm{\tau}_{i}=\left[\tau_{i}^{1},\ldots,\tau_{i}^{Q_{i}}\right]^{\top}\in\mathbb{R}^{Q_{i}}$ be the travel time observations for all trips on days $i$, where $Q_{i}$ is the total number of trips on day $i$. We define a transformation matrix $A_{i}=[a_{q,l}]\in\{0,1\}^{Q_{i}\times|\mathcal{V}|}$, where $a_{q,l}=1$ if trip $q$ uses link $l$ and 0 otherwise, and denote the $q$-th row in $A_{i}$ by $A_{i}^{q}$. We define $B_{i}=\operatorname{blkdiag}(\{A_{i}^{q}\})$ of size $Q_{i}\times Q_{i}|\mathcal{V}|$ as a block diagonal matrix composed of each row in $A_{i}$. Marginalizing the day-specific and trip-specific random effects in Eq. (2), the joint distribution of $\bm{\tau}_{i}$ can be derived as

### v2

@@ Abstract
Travel time estimation is a key task in navigation apps and web mapping services. Existing deterministic and probabilistic methods, based on the assumption of trip independence, predominantly focus on modeling individual trips while overlooking trip correlations. However, real-world conditions frequently introduce strong correlations between trips, influenced by external and internal factors such as weather and the tendencies of drivers. To address this, we propose a deep hierarchical joint probabilistic model ProbETA for travel time estimation, capturing both inter-trip and intra-trip correlations. The joint distribution of travel times across multiple trips is modeled as a low-rank multivariate Gaussian, parameterized by learnable link representations estimated using the empirical Bayes approach. We also introduce a data augmentation method based on trip sub-sampling, allowing for fine-grained gradient backpropagation when learning link representations. During inference, our model estimates the probability distribution of travel time for a queried trip, conditional on spatiotemporally adjacent completed trips. Evaluation on two real-world GPS trajectory datasets demonstrates that ProbETA outperforms state-of-the-art deterministic and probabilistic baselines, with Mean Absolute Percentage Error decreasing by over 12.60%. Moreover, the learned link representations align with the physical network geometry, potentially making them applicable for other tasks.
@@
Probabilistic travel time estimation, representation learning, uncertainty quantification, low-rank parameterization, distributional regression
@@ I Introduction
Travel Time Estimation (TTE) focuses on predicting the time it takes a vehicle to traverse a specific path connecting two locations (i.e., origin and destination). Given the departure time, TTE is equivalent to predicting the estimated time of arrival (ETA). TTE (or ETA prediction) is a fundamental application in navigation services and intelligent transportation systems: for example, individual travelers can better schedule their trips based on TTE, logistics, and delivery service operators can design more efficient and robust routes by taking TTE into consideration, and transport agencies can leverage TTE for better operation and management.
@@
ETA prediction has become an important problem in machine learning and has gained considerable attention in the literature. Most existing studies can be categorized into two groups based on whether path information is used to predict travel time [30, 1]. The first is the origin-destination-based (OD-based) approach (see, e.g., [28, 25, 12, 32, 14, 10, 26]). The key concept of this approach involves searching for past trips that have origin-destination profiles similar to the queried trip. The obtained travel time values are then used to calculate the mean and variance of the queried travel time [25]. Factors such as travel distance, departure time, road attributes, and other additional information are further introduced to enrich the feature space of OD pairs [12, 14, 26]. The periodic patterns of traffic conditions can also be analyzed to account for the temporal dynamics [32, 10, 26]. The OD-based method relies solely on structured features such as origin, destination, and time of day as covariates, and the models typically exhibit low computational complexity. Nonetheless, since travel time is primarily influenced by the specific route taken by a vehicle, the prediction accuracy of these OD-based models is often compromised due to the omission of path information. The presence of high-quality GPS data and advances in sequence modeling methodologies inspire researchers to use comprehensive trip information for TTE, resulting in the development of route-focused approaches. The key concept of the route-based approach involves learning representation vectors for road segments [8, 13] or GPS points[24] in a data-driven way and then combining these vectors with deep neural networks to obtain estimates. Most route-based methods utilize recursive neural networks to model temporal dynamics (see, e.g., [22, 31, 18, 29]) and graph neural networks to characterize spatial information (see, e.g., [9, 6, 27, 3]). In addition, attention mechanism [34, 2] and meta-learning [23, 5] are also incorporated into this category of methods. All the above works are deterministic models with an implicit assumption of trip independence, which neither provide confidence intervals for the estimates nor capture trip correlation, leading to significant limitations.
To quantify the uncertainty of the estimates, recent pioneering researches [33], [11] have focused on probabilistic TTE, which involves forecasting the probabilistic distribution of travel time on a specific path. These models generally assume that the travel time $\tau_{i}$ of a trip $i$ follows an independent Gaussian distribution $\tau_{i}\sim\mathcal{N}\left(f_{\mu}(X_{i}),f_{\sigma^{2}}(X_{i})\right)$, where $X_{i}$ is the input feature of trip $i$, and $f_{\mu}(\cdot)$ and $f_{\sigma^{2}}(\cdot)$ are two functions that produce the mean and variance of the distribution, respectively. In such models, the trip independence assumption is still being maintained, i.e., $Cov(\tau_{i},\tau_{j})=0,\forall i\neq j$. However, in real-world traffic scenarios, the travel times of two trips often exhibit strong correlation when they are influenced by the same factors (e.g., incidents and weather conditions). Those traditional models require incorporating all factors $F$ contributing to travel time correlation as conditional information (e.g., via feature embeddings) to satisfy the independence assumption of estimations, i.e. $p(\tau_{i},\tau_{j}|F_{\text{all}})=p(\tau_{i}|F_{\text{all}})p(\tau_{j}|F_{\text{all}})$. This is impractical, as external influencing factors are numerous, expensive, and difficult to model comprehensively. Disregarding any correlation factors can result in travel time being correlated, and makes their loss function under the independence assumption no longer valid[21, 19].
To address this problem, we aim to relax the independence assumption, directly model the joint probability distribution of trips, and explicitly learn the covariance matrix. However, this trip joint probability modeling approach is challenging due to: 1) the number of trip samples is large, and the maximum likelihood estimation based on the covariance matrix requiring $\mathcal{O}(N^{3})$ computational complexity; 2) Trips exhibit complex interaction, but the sparsity and coarseness of GPS trajectories make it challenging to capture and learn meaningful trip correlations accurately.
In this paper, we propose a novel deep hierarchical joint probabilistic model ProbETA that relaxes the independent distributed assumption to construct the joint distribution of multiple trips. It is a more general approach, optimized through maximum likelihood estimation of the joint distribution. Specifically, this model captures both inter-trip and intra-trip correlations by learning hierarchical random effects through learnable link representations. We employ a low-rank plus diagonal approach to construct the covariance matrix, effectively reducing the computational complexity of the model. Additionally, we introduce a trip sub-sampling data augmentation technique to efficiently leverage sparse GPS trajectories, enhancing correlation capture and link representation learning. The main contributions are summarized as follows:
@@
  • We propose a multi-trip joint probabilistic distribution model ProbETA for estimating travel times. Efficient low-rank parameterizations are introduced to learn inter-trip and intra-trip correlations among all segments in a transportation network.
@@
  • We propose a sub-sampling data augmentation approach to balance the samples and enhance the optimization effect in learning the link representation vectors. It facilitates specific optimization at the link level, enabling fine-grained modeling of link features.
@@
The remainder of this paper is organized as follows. Section II summarizes related work. Section III introduces key definitions and formulations. Section IV presents the key components of ProbETA, including parameterization of the inter-trip and intra-trip link covariance matrices and efficient likelihood evaluation. We evaluate the proposed model on two real-world datasets in Section V. Finally, we conclude this study in Section VI.
@@ II-A Deterministic Travel Time Estimation
ProbETA(ours) Route-based GPS Sequence Link Embedding Path-based-sum of Link Representation
@@ II-B Probabilistic Regression
In TTE, there are also some efforts to explore probabilistic modeling. Li et al. [11] proposed a deep generative model to learn the travel time distribution by conditioning on the real-time traffic. This model interprets the generation of travel time using a three-layer hierarchical probabilistic model, which captures dynamically changing real-time traffic conditions and static spatial features, and then generates estimation times based on an attention mechanism. In [33], Zhou et al. proposed to learn the local representations of road segments over a temporal attributed graph, by jointly exploiting the dynamic traffic conditions and the topology of the road networks. Then a distribution loss based on the negative log-likelihood (NLL) is developed to fulfill the purpose of travel time distribution estimation. Liu et al. [15] transforms the travel time estimation task into a multi-class classification problem to account for travel time uncertainty and introduces an adaptive local label-smoothing scheme to capture ordinal relationships between labels. Mao et al. [16] introduced GMDNet, a Graph-based Mixture Density Network, which harnesses the advantages of both graph neural networks and mixture density networks for estimating travel time distribution. They utilized the Expectation-Maximization (EM) framework to enhance stability during training. In addition, tensor-based methods have also been used to construct probabilistic estimation models for travel time.
@@
Figure 1: Overall architecture of ProbETA. The black flow represents the shared process between training and inference, the green flow represents the training process, and the blue flow represents the inference process.
@@ III-B Definition: Links and Trips
A link $l\in\mathcal{V}$ refers to a road segment. Each link in the road network has a unique index. In this work, we follow OpenStreetMap (OSM) [7] to define links. The trajectory of a trip is an ordered sequence of time-stamped GPS records. For a trip with $k$ GPS data points, we denote its trajectory by $T_{g}=\left\{(l_{1},c_{1}),\ldots,(l_{k},c_{k})\right\}$, where the time index $c$ is monotonically increasing. The total travel time of the trip is $\tau=c_{k}-c_{1}$. Note that it is possible that (1) more than one GPS data points are located on the same link, and (2) a traversed link is not captured if no GPS records are registered. The link index sequence, after completing the entire sequence and removing duplicates, is used as model input, denoted by ${T}=\{l_{1},...,l_{k}\}$. In other words, a trip $T$ can be regarded as an ordered set of traveled links, with all relevant information derived from the raw GPS sequence.
@@ IV Methodology
IV-A Overview of ProbETA
The overall architecture of ProbETA is shown in Figure 1. A unique property of ProbETA is that the correlation among multiple trips is explicitly modeled, and the travel times for multiple trips are jointly modeled as a multivariate Gaussian distribution. Each trip is represented by aggregating the low-rank representation of all those links that it covers. A neural network-based mapping function is constructed to learn the mean travel time based on the trip representation vectors, while the covariance is formed using the inner product of these trip representation vectors for simultaneously modeling inter- and intra-trip correlations. Additionally, trip sub-sampling is used to augment the training data, balancing the samples and allowing gradients to differentially impact the representation at the link level within the same trip. Finally, historical trips in spatiotemporal proximity to the query trip are introduced to obtain the travel time distribution of a queried trip conditional on those observed trips. The final estimations are obtained by randomly sampling the conditional distribution. We explain the details of each component in the following subsections.
@@ IV-B Parameterizing Link Travel Time Distribution
In this study, we focus on modeling trip travel time over a specific time horizon (e.g., 8:00–9:00 AM) of weekdays. In doing so, we restrict our scope to only model the variations in link travel time resulting from external factors rather than traffic demand. Given that the historical data spans several days, we employ a hierarchical model to describe both daily random fluctuations and random effects at the trip level. In particular, we construct a three-tier hierarchical model for link travel time model the travel time $t_{l,i,q}$ for link $l$, day $i$, and trip $q$ as
@@
where $\mu_{l}$ represents the overall mean travel time for link $l$ across all days, which can be produced by learned link representation. $\eta_{l,i}$ accounts for day-specific deviations as a result of the impacts of unobserved factors that affects all trips in a day but are not used in modeling $\mu_{l}$, such as weather conditions, road works and public holidays; and $\epsilon_{l,i,q}$ captures the trip-specific error (e.g., differences in driver/vehicle profile, or short-duration delays due to traffic incidents affecting multiple nearby links). Note that a day can be divided into multiple time periods, and deploying our model within each period can yield more effective estimations.
@@
$\Sigma_{d}=\varphi_{\Sigma_{d}}(L)\varphi_{\Sigma_{d}}(L)^{\top}=L\bm{w}_{\Sigma_{d}}\bm{w}_{\Sigma_{d}}^{\top}L^{\top},$ (4)
where $L\in\mathbb{R}^{|\mathcal{V}|\times r_{L}}$ and $r_{L}\ll|\mathcal{V}|$. With this assumption, $\Sigma_{d}$ is positive semi-definite and we can consider each row of $L$ as a representation vector of a link, $\varphi_{\Sigma_{d}}(\cdot)$ is a branching function composed of a multilayer perception with learnable parameters $\bm{w}_{\Sigma_{d}}\in\mathbb{R}^{r_{L}\times r_{L}}$. Then, we adopt a regression function $f_{\mu}(\cdot)$ to learn the global mean of travel time based on the other branch of representation matrix $\varphi_{\mu}(L)$
$\bm{\mu}=f_{\mu}(\varphi_{\mu}(L))=L\bm{w}_{\mu}\bm{r}_{\mu},$ (5)
where $\bm{w}_{\mu}\in\mathbb{R}^{r_{L}\times r_{L}}$, $\bm{r}_{\mu}\in\mathbb{R}^{r_{L}}$ is a parameter vector to be learned. Combining Eqs. (5) and (4), we can specify the joint Gaussian distribution based on the link representation matrix $L$. It should be noted that the modeling of $\bm{\mu}$ is not limited to the segment feature $L$. One can also introduce other features that are available, such as driver/vehicle profile, time of day, day of week, public holiday information, and weather conditions. This will create a more comprehensive mean process. For example, Ref. [17] introduces representation learning for rare temporal conditions such as events and holidays.
Figure 2: Graph node contraction. Aggregating multiple nodes into new nodes, this contraction property can be used to simplify the graph structure or construct trip representations.
In addition to computational efficiency, the proposed specification also provides excellent contraction properties and facilitates the aggregation of unimportant nodes in $\mathcal{V}$. For example, consider $\{2,9,10\}$ as unimportant nodes, the contractions of links $\{1,2\}$ and $\{8,9,10\}$ in Figure 2 can be achieved by introducing a $7\times 10$ binary mapping matrix
@@
Given the affine property of the Gaussian distribution, the mean travel time in the simplified network is $\bm{\mu}^{\prime}=M\bm{\mu}=ML\bm{w}_{\mu}$ and the covariance matrix becomes $\Sigma_{d}^{\prime}=ML(ML)^{\top}$. This corresponds to having a new representation matrix $L^{\prime}=ML$. This affine property can further help us in deriving the distribution of trip travel time by aggregating the traversed links. We define the representation of a trip (as a row vector) to be the path-based-sum of link representation vectors, i.e., $ML$, with $M\in\{0,1\}^{1\times|\mathcal{V}|}$. The mean travel time on a path becomes $M(L\bm{w}_{\mu})=(ML)\bm{w}_{\mu}$ can be directly computed from the trip representation $ML$.
@@
$\Sigma_{p}=\varphi_{\Sigma_{p}}(H)\varphi_{\Sigma_{p}}(H)^{\top}+D=H\bm{w}_{\Sigma_{p}}\bm{w}_{\Sigma_{p}}^{\top}H^{\top}+D,$ (7)
where $H\in\mathbb{R}^{|\mathcal{V}|\times r_{H}}$ with $r_{H}\ll|\mathcal{V}|$ and $D\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ is a diagonal matrix with $D_{i,i}>0$ for $i=1,\ldots,|\mathcal{V}|$, $\varphi_{\Sigma_{p}}(\cdot)$ is a branching function composed of a multilayer perception with learnabel parameters $\bm{w}_{\Sigma_{p}}\in\mathbb{R}^{r_{H}\times r_{H}}$. We use a Softplus function nested with a regression function $f_{d}(\cdot)$ to model the diagonal entries (noise variance) $\text{diag}(D)=\log\left(1+\exp(f_{d}(\varphi_{d}(H)))\right)$ where $\varphi_{d}(H)=H\bm{w}_{d}$ is a branching function and $\bm{w}_{d}\in\mathbb{R}^{r_{H}\times r_{H}}$. Similar to $L$, we can also perform contractions on the learned matrix $H$ and $\operatorname{diag}(D)$.
IV-C Joint Distribution for Multiple Trips
A key challenge in learning the link representations is that trip-level link travel time $t_{l,i,q}$ is not observable in the raw GPS data. Instead, what we have is the final travel time for a trip. Let $\bm{\tau}_{i}=\left[\tau_{i}^{1},\ldots,\tau_{i}^{Q_{i}}\right]^{\top}\in\mathbb{R}^{Q_{i}}$ be the travel time observations for all trips, where $Q_{i}$ is the total number of trips on day $i$. We define a transformation matrix (or a trip-link incident matrix) $A_{i}=[a_{q,l}]\in\{0,1\}^{Q_{i}\times|\mathcal{V}|}$, where $a_{q,l}=1$ if trip $q$ uses link $l$ and 0 otherwise, and denote the $q$-th row in $A_{i}$ by $A_{i}^{q}$. We define $B_{i}=\operatorname{blkdiag}(\{A_{i}^{q}\})$ of size $Q_{i}\times Q_{i}|\mathcal{V}|$ as a block diagonal matrix composed of each row in $A_{i}$. Marginalizing the day-specific and trip-specific random effects in Eq. (2), the joint distribution of $\bm{\tau}_{i}$ is derived as